OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. 4x + 4y = 6 x + 4y = 12 4x + 4y = –12 x – 4y = –6

OpenStudy (anonymous):

slope = (3-2)/(6-2) = 1/4. So, y=x/4 + b When x = 2 then y = 2: 2 = 1/2 + b => b= 3/2. So, y = x/4 + 3/2 4y = x + 6 4y-x =6 or x - 4y = -6 is the answer